OpenStudy (anonymous):

PLEASE HELP!!! in the simplest a+bi form, the roots of the equation 2+(5/x^2)=6/x

OpenStudy (anonymous):

Strange. I haven't done this in a while... I might refer you to @thomaster for help. :) This stuff really isn't my strong point. Sorry about that! :)

OpenStudy (anonymous):

its alright!! thank you so much though!! xo

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

I don't know how "standard" my method is, but this is the way I'd do it. To go ahead and make the equation simpler, multiply both sides by x.

OpenStudy (anonymous):

or, x^2 to eliminate ALL denominators.

OpenStudy (anonymous):

Quite.

OpenStudy (anonymous):

i did the multiplying by x^2 on both side and the answer came out to be x=3 now how do i write in a+bi form? is that right?

OpenStudy (triciaal):

2 + 5/x^2 = 6/x multiply by x^2 and rewrite to get 2x^2-6x +5 =0 a=2 b=-6 c=5 use quadratic formula then rt-4 express as rt-1.rt4

OpenStudy (anonymous):

@triciaal so what will be the answer in a+bi form?

OpenStudy (triciaal):

what do u have right now

OpenStudy (triciaal):

-3/2 +1/2 i and

OpenStudy (anonymous):

Alright! thank you soo much!! xo